OpenStudy (anonymous):

1. at my birthday party, the girls ate 31/2 pizzas and the boys at 7 1/2. how many pizzas were eaten at my party? 2. i bought 2 1/2 gallons of paint but i only used 2/4 gallons of the paint. How much paint do i have left? 3. my recipe calls for 2/3 cups of white flour and 2 1/5 cups of whole wheat flour. how much flour do i need in total for my recipe? 4. My dog is 5 1/2 years old. My cat is 4 1/4 years younger than my dog. How old is my cat? 5. During the pie eating contest my dad ate 5 3/4 pies and my mom ate 2 1/4 pies. How many pies did they eat altogethe

OpenStudy (akshay_budhkar):

OMG!!! you typed this!! lol

OpenStudy (anonymous):

Is that your whole textbook?

OpenStudy (akshay_budhkar):

lol imran

OpenStudy (anonymous):

O___O

OpenStudy (anonymous):

hey srry ik its messy

OpenStudy (anonymous):

lemme break this down, see if that helps: 1) 31/2 + 7 1/2= 2) 2 1/2 - 2/4= 3) 2/3 + 2 1/5= 4) 5 1/2 - 4 1/4= 5) 5 3/4 + 2 1/4= -- also in this scenario your parents are now diabetic

OpenStudy (anonymous):

kk